# Apparent delay of the Kibble-Zurek mechanism in quenched open systems

### v1

@@ III.1 Signatures of the delayed KZM
Figure 4: (a-b) Sketch of the (a) open Dicke model and the (b) open Dicke lattice model with periodic boundary condition. (c-e) Exemplary dynamics of the (c) occupation number, (d) phase of the $\hat{a}$-mode, and (e) $S^{x}$ of the quenched open Dicke model for $\omega\tau_{q}=1000$ and $\varepsilon_{f}=0.2$. (f-h) Exemplary steady-state spatial distribution of the (f) occupation number, (g) phase of the $\hat{a}_{\ell}$ mode, and (h) $S^{x}_{\ell}$ of the open Dicke lattice model in the SR phase. (i-j) Signatures of the delayed KZM for the (i) open DM and (j) open DLM. (Top Panel) Scaling of $\hat{t}_{\mathrm{th}}$ as a function of $\tau_{q}$ for $\kappa=0.1\omega$. The vertical dashed lines represent $\tau_{q,c}$, while the solid lines corresponds to $\tau_{q,c}^{*}$. (Bottom Panel) The boundary of the delayed KZM regime as a function of $\kappa$. The remaining parameters are set to $\varepsilon_{f}=0.2$ and $|a|^{2}_{\mathrm{th}}=2$ for the open DM, and $J=0.1\omega$, $\varepsilon_{f}=0.25$, and $|a|^{2}_{\mathrm{th}}/M=2$, with $M=500$ sites, for the open DLM.
We now test whether the delayed KZM is a generic feature of open systems by considering two fully-connected systems: the open DM, schematically represented in Fig. 4(a), and its one-dimensional lattice version, the open DLM, as shown in Fig. 4(b). Both systems are described by the master equation in Eq. (2), with the Hamiltonian of the open DM being [41, 42]
@@
while the Hamiltonian of its $M$-site lattice version with periodic boundary condition takes the form [43]
@@
The open DM has the same dissipator as the open COS, while the dissipator of the open DM is $\mathcal{D}\hat{\rho}=\kappa\sum_{\ell}^{M}\left(2\hat{a}_{\ell}\hat{\rho}\hat{a}_{\ell}^{\dagger}-\left\{\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell},\hat{\rho}\right\}\right)$ [43]. The open DM describes the dynamics of $N$ two-level systems, represented by the collective spin operators $\hat{S}^{x,y,z}$, coupled to a dissipative bosonic mode, $\hat{a}$, which in cavity-QED experiments corresponds to a photonic mode [27, 40, 42, 47]. In both systems, $\omega$ and $\omega_{0}$ are the bosonic and spin transition frequencies, respectively, and $\lambda$ is the spin-boson coupling, while $J$ represents the nearest-neighbour interaction in the open DLM.
@@
In equilibrium, the open DM has two phases: the NP and the superradiant phase (SR) [41, 42]. The NP is characterized by a fully polarised collective spin at the $-z$ direction, i.e. $S^{z}=-N/2$, and a zero total occupation number, $|a|^{2}$. Meanwhile, the SR phase is associated with the $\mathbb{Z}_{2}$ symmetry breaking of the system, leading to a nonzero $S^{x}$ and $|a|^{2}$, with $S^{x}$ ($a$) picking a random sign (phase) from the two degenerate steady states of the system [42]. The two phases are separated by the same critical point as the open COS [42]. Under a finite quench, however, the open DM exhibits non-trivial dynamics as it transitions from the NP to the SR phase. We present in Figs. 4(c) to 4(e) an exemplary dynamics of the total occupation number, $|a|^{2}$, and the phase of the bosonic mode, $\varphi$, and $S^{x}$ of the open DM for $\omega\tau_{q}=1000$. We initialized the system near the steady state of the NP, where the initial values of the bosonic mode are $a=0.01$, while the collective spin operators are
$\quad S^{x}(t_{i})=\frac{N}{2}\delta,\quad S^{y}(t_{i})=0,\quad S^{z}(t_{i})=\frac{N}{2}\sqrt{1-\delta^{2}},$ (12)
@@
where $\delta$ is a perturbation set to $\delta=0.01$. We can observe that when the system is in the NP, the occupation number approaches the NP steady state, $a=0$, which is consistent with our predicted behaviour from the potential surface interpretation of phase transition in an open system, which we describe in Sec. II. In addition, the phase of the bosonic mode oscillates from $-\pi$ to $\pi$, while the $S^{x}$ remains close to $S^{x}=0$. As $\varepsilon(t)>0$ at $t>0$, the system enters the SR phase, which results in $\varphi$ spontaneously admitting a finite value as the $|a|^{2}$ starts to exponentially grow until $t=\varepsilon_{f}\tau_{q}$, where the ramp terminates. At that point, the $|a|^{2}$ finally saturates at the steady state of the SR phase. Meanwhile, the transition of $S^{x}$ from its behaviour in the NP to the SR phase only becomes prominent at a later time. We will further expand on the implication of the behaviour of these order parameters later in Sec. IV.
@@
As for the open DLM, for small values of $J$, the interaction between the open DMs modifies the $\lambda_{c}$ into a critical line [43],
@@
Moreover, suppose that we drive the open DLM from the NP to the SR phase using a finite quench after initializing it near the steady state of NP. Specifically, we initialize the collective spins at $S^{x,y,z}_{\ell}=S^{x,y,z}(t=t_{i})$, while the bosonic modes are initialized at the vacuum state, which can be represented as a complex Gaussian variable $a_{\ell}=\frac{1}{2}\left(\eta_{\ell}^{\mathrm{R}}+\eta_{\ell}^{\mathrm{I}}\right)$, where $\eta_{\ell}^{R,I}$ are random numbers sampled from a Gaussian distribution satisfying $\eta_{\ell}^{i}\right>=0$ and $\eta_{\ell}^{i}\eta_{m}^{j}\right>=\delta_{i,j}\delta_{\ell,m}$ for $i,j=R,I$ [51]. Then, as the system enters the SR phase, each site can independently pick between the two degenerate steady states available, allowing for the formation of domains and point defects, the number of which depends on the correlation length of the system. We present in Figs. 4(f) to 4(h) the exemplary spatio-temporal dynamics of $|a_{\ell}|^{2}$, $\varphi_{\ell}$, and $S^{x}_{\ell}$ of the open DLM after doing a finite quench towards the SR phase. We can observe that the point defects can manifest either as dips in the occupation number, phase slips in the spatial profile of $\varphi_{\ell}$, or domain walls in $S^{x}_{\ell}$. Note that the defect number $N_{d}$ follows the predicted KZM power-law scaling with $\tau_{q}$, which we demonstrate in Appendix B. Since the notion of topological defects is well-defined in the open DLM, it serves as a good test bed for the delayed KZM for systems with short-range interaction. This is in addition to the open DM, which has been experimentally shown to exhibit signatures of the KZM [27] despite the open question of its non-equilibrium universality class [52, 53].
@@
$\frac{\hat{H}^{\mathrm{DLM}}}{\hbar}\approx\frac{1}{\hbar}\sum_{k}\hat{H}^{\mathrm{OM}}_{k}$ (17)
@@
is the Hamiltonian of each uncoupled oscillator at the momentum mode $k$ and $\omega_{k}=\omega-2J\cos(k)$. In this form, we can easily observe that the open DLM has a similar structure to the COS, with the similarity being more apparent at the zero-momentum mode,
$\frac{\hat{H}^{\mathrm{OM}}_{0}}{\hbar}=\left(\omega-2J\right)\hat{a}_{0}^{\dagger}\hat{a}_{0}+\omega_{0}\hat{b}_{0}^{\dagger}\hat{b}_{0}+\lambda\left(\hat{a}_{0}^{\dagger}+\hat{a}_{0}\right)\left(\hat{b}_{0}^{\dagger}+\hat{b}_{0}\right)$ (19)
@@
Since we have shown the generality of the delayed KZM on open systems, we now explore in greater detail the dissipative and near-critical nature of the delayed KZM on Sec. III.2.
@@ III.2 Dissipative and critical nature of the delayed KZM
Figure 5: (a) Dependence of the occupation number decay rate of the open DLM on $\tau_{q}$ for $\varepsilon_{f}=0.25$ and $\kappa=J=0.1\omega$. The solid line corresponds to the average value of $\gamma_{d}$ from $\omega\tau_{q}=10^{2}$ to $\omega\tau_{q}=10^{4}$. (b-c) $\bar{\gamma}_{d}$ as a function of (b) $\varepsilon_{f}$ for $\kappa=0.1\omega$ and (c) $\kappa$ for $\varepsilon_{f}=0.25$. The dashed lines correspond to the best-fit lines.
@@
Figure 6: (a-b) Scaling of (a) $\hat{t}_{\mathrm{th}}$ and (b) $\varepsilon(\hat{t}_{\mathrm{th}})$ of the open DLM as a function of $\tau_{q}$ for different $\varepsilon_{f}$. The dashed lines in (a) correspond to the best-fit lines of $\hat{t}_{\mathrm{th}}$ for fast and slow quenches, with their intersections marking $\tau_{q,c}$. Meanwhile, the dashed lines in (b) denote $\tau_{q,c}^{*}$. (c) Borders of the delayed KZM regime as a function of $\varepsilon_{f}$. The dashed line corresponds to the best-fit line of $\tau_{q,c}^{*}$. The remaining parameters are $\kappa=J=0.1\omega$, and $|a|^{2}_{\mathrm{th}}/M=2$, with $M=500$ sites.
Given that $\varepsilon_{f}$ do not alter the behaviour of the decay rate of $|a|^{2}$, it is natural to ask whether varying $\varepsilon_{f}$ has any significant effect as well on the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$, and on the signatures of the delayed KZM. We answer the first question in Figs. 6(a) and 6(b), where we show the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$, respectively, with $\tau_{q}$ for different values of $\varepsilon_{f}$. We can see that varying $\varepsilon_{f}$ does not significantly change the scaling of the KZM quantities considered. However, the $\tau_{q,c}^{*}$, shown as solid lines in Fig. 6(b), increases significantly as we decrease $\varepsilon_{f}$. This modification on $\tau_{q,c}^{*}$ becomes more apparent in Fig. 6(c), where we show the scaling of $\tau_{q,c}$ and $\tau_{q,c}^{*}$ as a function of $\varepsilon_{f}$. Notice that both quantities are inversely proportional to $\varepsilon_{f}$, with $\tau_{q,c}^{*}$ dropping faster than $\tau_{q,c}$ as $\varepsilon\rightarrow\infty$. As a result, the delayed KZM regime vanishes for large $\varepsilon_{f}$, highlighting that its signatures become more apparent for strongly dissipative systems quenched near criticality. We finally note that $\tau_{q,c}^{*}$ follows a power-law scaling as evidenced by the power-law fit curve shown in Fig. 6(c). In particular, since $\tau_{q,c}^{*}$ becomes the true critical quench time separating the breakdown and validity of the KZM at large $\varepsilon_{f}$, we expect that it should follow the power-law scaling [39]
@@
So far, we have shown that the presence of the delayed KZM leads to a significant deviation between the true freeze-out time, $\hat{t}$ and the transition time, $\hat{t}_{\mathrm{th}}$. Given that the delayed KZM becomes more prominent near criticality at strong dissipation, we now address in the next section how the threshold-based criterion for determining $\hat{t}_{\mathrm{th}}$ contributes to the deviation and whether a more accurate method can be used to measure $\hat{t}$.
@@ IV Transition Time Measurement
Figure 7: (a-b) Scaling of (a) $\hat{t}_{\mathrm{th}}$ and (b) $\varepsilon(\hat{t}_{\mathrm{th}})$ of the open DM as a function of $\tau_{q}$ for different values of $|a|^{2}_{\mathrm{th}}$. (c) Boundary of the delayed KZM regime as a function of $\kappa$ for different $|a|^{2}$. The remaining parameters are set to $\varepsilon_{f}=0.2$ and $\kappa=0.1\omega$.
We answer the first question in Figs. 7(a) and 7(b) where we present the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$, respectively. For this part, while we only consider the open DM, the results here should apply to the COS and the open DLM as well. We can observe that the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ do not significantly change as we increase the threshold value. In particular, while the $\hat{t}_{\mathrm{th}}$ is only shifted by a constant value as $|a|^{2}_{\mathrm{th}}$ increases, both KZM quantities considered eventually collapses in a single scaling as $\tau_{q}\rightarrow\infty$. As for the boundaries of the delayed KZM regime, we can observe in Fig. 7(c) that the gap between $\tau_{q,c}$ and $\tau_{q,c}^{*}$ widens as we increase $|a|^{2}_{\mathrm{th}}$, implying that the delayed KZM becomes more prominent at large $|a|^{2}_{\mathrm{th}}$.
We can understand the widening of the delayed KZM regime for large $|a|^{2}_{\mathrm{th}}$ by noting that in an ideal set-up where $\hat{t}$ can be accurately identified, the gap between $\tau_{q,c}$ and $\tau_{q,c}^{*}$ vanishes, and thus following the prediction in Ref. [39], $\tau_{q,c}=\tau_{q,c}^{*}=\hat{t}/\varepsilon_{f}$. Since for any threshold-based criterion, $\tau_{q,c}^{*}=\hat{t}_{\mathrm{th}}/\varepsilon_{f}$ and $\tau_{q,c}\neq\tau_{q,c}^{*}$ for large $\kappa$ and small $\varepsilon_{f}$, then
@@
Let us assume that within the time interval $[\hat{t},\epsilon_{f}\tau_{q}]$, the total occupation is exponentially growing such that $|a|^{2}\propto\exp(\gamma_{g}t)$, where $\gamma_{g}$ is the growth rate of the total occupation number. This assumption is supported by Fig. 4(c), where the $|a|^{2}$ of the open DM exponentially grows from the minimum value to its saturation value. With this assumption we can infer that $\hat{t}_{\mathrm{th}}\propto\ln|a|^{2}_{\mathrm{th}}/\gamma_{g}$ and $\hat{t}\propto\ln|a|^{2}_{\mathrm{min}}/\gamma_{g}$, where $|a|^{2}_{\mathrm{min}}$ is the minimum value of the total occupation number. Thus,
@@
Figure 8: (a) Ramp protocol for the quench time within the $\tau_{q}$-regime of the delayed KZM for the open DLM. The circle point marks $\varepsilon(\hat{t})$, while the diamond point denotes $\varepsilon(\hat{t}_{\mathrm{th}})$. (b-c) Spatio-temporal dynamics of (b) the occupation number $|a_{\ell}|^{2}$ and (c) the phase of the $\hat{a}_{\ell}$-mode, $\varphi_{\ell}$. (d) Exemplary dynamics of the defect number and the site-averaged $\varphi_{\ell}$. The vertical dashed lines correspond to $\omega\hat{t}$, while the solid lines represent $\omega\hat{t}_{\mathrm{th}}$. The remaining parameters are set to $\varepsilon_{f}=1.25$, $\kappa=J=0.1\omega$, and $|a|^{2}_{\mathrm{th}}=0.05|a|^{2}_{s}$, where $|a|^{2}_{s}$ is the steady state of $|a|^{2}$.
Now, determining an optimal threshold value that suppresses the signatures of the delayed KZM may be difficult to achieve as it requires prior knowledge of $|a|^{2}_{\mathrm{min}}$ for arbitrary $\tau_{q}$. This problem motivates the question of whether an alternative method can be used to infer $\hat{t}$ without relying on any threshold-based criterion. As we have hinted in the dynamics of the phase of the $\hat{a}$-mode of the open DM shown in Fig. 4(d), we can do this by choosing an appropriate order parameter that rapidly reaches its steady state upon the system entering a phase transition. In the case of the open DM, this order parameter corresponds to the boson mode’s phase, $\varphi$. We demonstrate this method further in Figs. 8(b) and 8(c), where we show that for non-zero dimensional systems, like the open DLM, we can use the phase information of the $\hat{a}_{\ell}$-modes to extract $\hat{t}$. As presented in Fig. 8(d), we can do this by determining the time at which either the defect number, $N_{d}$, or the site-averaged phase begins to saturate. In the COS level, the inferred $\hat{t}$ for this method would be equivalent to the moment the system picks a new global minimum it would fall onto, signalling phase transition. Thus, we expect that if the system’s phase information is available in an experimental set-up, such as in Ref. [54], then that can serve as a more sensitive tool for detecting phase transitions compared to threshold-based order parameters that depend on the mode occupations.

### v2

@@ III.1 Signatures of the delayed KZM
Figure 4: (a), (b) Sketch of the (a) open Dicke model and the (b) open Dicke lattice model with periodic boundary condition. (c)-(e) Exemplary dynamics of the (c) occupation number, (d) phase of the $\hat{a}$ mode, and (e) $S^{x}$ of the quenched open Dicke model for $\omega\tau_{q}=1000$ and $\varepsilon_{f}=0.2$. (f)-(h) Exemplary steady-state spatial distribution of the (f) occupation number, (g) phase of the $\hat{a}_{\ell}$ mode, and (h) $S^{x}_{\ell}$ of the open Dicke lattice model in the SR phase. (i)-(j) Signatures of the delayed KZM for the (i) open DM and (j) open DLM. Top panel: Scaling of $\hat{t}_{\mathrm{th}}$ as a function of $\tau_{q}$ for $\kappa=0.1\omega$. The vertical dashed lines represent $\tau_{q,c}$, while the solid lines corresponds to $\tau_{q,c}^{*}$. Bottom panel: The boundary of the delayed KZM regime as a function of $\kappa$. The remaining parameters are set to $\varepsilon_{f}=0.2$ and $|a|^{2}_{\mathrm{th}}=2$ for the open DM, and $J=0.1\omega$, $\varepsilon_{f}=0.25$, and $|a|^{2}_{\mathrm{th}}/M=2$, with $M=500$ sites, for the open DLM.
We now test whether the delayed KZM is a generic feature of open systems by considering two fully connected systems: the open DM, schematically represented in Fig. 4(a), and its one-dimensional lattice version, the open DLM, as shown in Fig. 4(b). Both systems are described by the master equation in Eq. (2), with the Hamiltonian of the open DM being [42, 43]
@@
while the Hamiltonian of its $M$-site lattice version with periodic boundary conditions takes the form [44]
@@
The open DM has the same dissipator as the open COS, while the dissipator of the open DM is $\mathcal{D}\hat{\rho}=\kappa\sum_{\ell}^{M}\left(2\hat{a}_{\ell}\hat{\rho}\hat{a}_{\ell}^{\dagger}-\left\{\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell},\hat{\rho}\right\}\right)$ [44]. The open DM describes the dynamics of $N$ two-level systems, represented by the collective spin operators $\hat{S}^{x,y,z}$, coupled to a dissipative bosonic mode, $\hat{a}$, which in cavity-QED experiments corresponds to a photonic mode [28, 41, 43, 48]. In both systems, $\omega$ and $\omega_{0}$ are the bosonic and spin transition frequencies, respectively, and $\lambda$ is the spin-boson coupling, while $J$ represents the nearest-neighbor interaction in the open DLM.
@@
In equilibrium, the open DM has two phases: the NP and the superradiant (SR) phase [42, 43]. The NP is characterized by a fully polarised collective spin at the $-z$ direction, i.e. $S^{z}=-N/2$, and a zero total occupation number, $|a|^{2}$. Meanwhile, the SR phase is associated with the $\mathbb{Z}_{2}$ symmetry breaking of the system, leading to a nonzero $S^{x}$ and $|a|^{2}$, with $S^{x}$ ($a$) picking a random sign (phase) from the two degenerate steady states of the system [43]. The two phases are separated by the same critical point as the open COS [43]. Under a finite quench, however, the open DM exhibits nontrivial dynamics as it transitions from the NP to the SR phase. We present in Figs. 4(c)-4(e) an exemplary dynamics of the total occupation number, $|a|^{2}$, and the phase of the bosonic mode, $\varphi$, and $S^{x}$ of the open DM for $\omega\tau_{q}=1000$. We initialized the system near the steady state of the NP, where the initial values of the bosonic mode are $a=0.01$, while the collective spin operators are
$\quad S^{x}(t_{i})=\frac{N}{2}\delta,\quad S^{y}(t_{i})=0,\quad S^{z}(t_{i})=-\frac{N}{2}\sqrt{1-\delta^{2}},$ (12)
@@
where $\delta$ is a perturbation set to $\delta=0.01$. We can observe that when the system is in the NP, the occupation number approaches the NP steady state, $a=0$, which is consistent with our predicted behavior from the potential surface interpretation of phase transition in an open system, which we describe in Sec. II. In addition, the phase of the bosonic mode oscillates from $-\pi$ to $\pi$, while the $S^{x}$ remains close to $S^{x}=0$. As $\varepsilon(t)>0$ at $t>0$, the system enters the SR phase, which results in $\varphi$ spontaneously admitting a finite value as the $|a|^{2}$ starts to exponentially grow until $t=\varepsilon_{f}\tau_{q}$, where the ramp terminates. At that point, the $|a|^{2}$ finally saturates at the steady state of the SR phase. Meanwhile, the transition of $S^{x}$ from its behavior in the NP to the SR phase only becomes prominent at a later time. We will further expand on the implication of the behavior of these order parameters later in Sec. IV.
@@
As for the open DLM, for small values of $J$, the interaction between the open DMs modifies the $\lambda_{c}$ into a critical line [44]:
@@
Moreover, suppose that we drive the open DLM from the NP to the SR phase using a finite quench after initializing it near the steady state of NP. Specifically, we initialize the collective spins at $S^{x,y,z}_{\ell}=S^{x,y,z}(t=t_{i})$, while the bosonic modes are initialized at the vacuum state, which can be represented as a complex Gaussian variable $a_{\ell}=\frac{1}{2}\left(\eta_{\ell}^{\mathrm{R}}+\eta_{\ell}^{\mathrm{I}}\right)$, where $\eta_{\ell}^{R,I}$ are random numbers sampled from a Gaussian distribution satisfying $\eta_{\ell}^{i}\right>=0$ and $\eta_{\ell}^{i}\eta_{m}^{j}\right>=\delta_{i,j}\delta_{\ell,m}$ for $i,j=R,I$ [52]. Then, as the system enters the SR phase, each site can independently pick between the two degenerate steady states available, allowing for the formation of domains and point defects, the number of which depends on the correlation length of the system. We present in Figs. 4(f)-4(h) the exemplary spatiotemporal dynamics of $|a_{\ell}|^{2}$, $\varphi_{\ell}$, and $S^{x}_{\ell}$ of the open DLM after doing a finite quench towards the SR phase. We can observe that the point defects can manifest either as dips in the occupation number, phase slips in the spatial profile of $\varphi_{\ell}$, or domain walls in $S^{x}_{\ell}$. Note that the defect number $N_{d}$ follows the predicted KZM power-law scaling with $\tau_{q}$, which we demonstrate in Appendix B. Since the notion of topological defects is well-defined in the open DLM, it serves as a good test bed for the delayed KZM for systems with short-range interaction. This is in addition to the open DM, which has been experimentally shown to exhibit signatures of the KZM [28] despite the open question of its nonequilibrium universality class [53, 54].
@@
$\frac{\hat{H}^{\mathrm{DLM}}}{\hbar}\approx\frac{1}{\hbar}\sum_{k}\hat{H}^{\mathrm{OM}}_{k},$ (17)
@@
is the Hamiltonian of each uncoupled oscillator at the momentum mode $k$ and $\omega_{k}=\omega-2J\cos(k)$. In this form, we can easily observe that the open DLM has a similar structure to the COS, with the similarity being more apparent at the zero-momentum mode:
$\frac{\hat{H}^{\mathrm{OM}}_{0}}{\hbar}=\left(\omega-2J\right)\hat{a}_{0}^{\dagger}\hat{a}_{0}+\omega_{0}\hat{b}_{0}^{\dagger}\hat{b}_{0}+\lambda\left(\hat{a}_{0}^{\dagger}+\hat{a}_{0}\right)\left(\hat{b}_{0}^{\dagger}+\hat{b}_{0}\right).$ (19)
@@
Since we have shown the generality of the delayed KZM on open systems, we now explore in greater detail the dissipative and near-critical nature of the delayed KZM in Sec. III.2.
@@ III.2 Dissipative and critical nature of the delayed KZM
Figure 5: (a) Dependence of the occupation number decay rate of the open DLM on $\tau_{q}$ for $\varepsilon_{f}=0.25$ and $\kappa=J=0.1\omega$. The solid line corresponds to the average value of $\gamma_{d}$ from $\omega\tau_{q}=10^{2}$ to $\omega\tau_{q}=10^{4}$. (b), (c) $\bar{\gamma}_{d}$ as a function of (b) $\varepsilon_{f}$ for $\kappa=0.1\omega$ and (c) $\kappa$ for $\varepsilon_{f}=0.25$. The dashed lines correspond to the best-fit lines.
@@
Figure 6: (a)-(b) Scaling of (a) $\hat{t}_{\mathrm{th}}$ and (b) $\varepsilon(\hat{t}_{\mathrm{th}})$ of the open DLM as a function of $\tau_{q}$ for different $\varepsilon_{f}$. The dashed lines in (a) correspond to the best-fit lines of $\hat{t}_{\mathrm{th}}$ for fast and slow quenches, with their intersections marking $\tau_{q,c}$. Meanwhile, the dashed lines in (b) denote $\tau_{q,c}^{*}$. (c) Borders of the delayed KZM regime as a function of $\varepsilon_{f}$. The dashed line corresponds to the best-fit line of $\tau_{q,c}^{*}$. The remaining parameters are $\kappa=J=0.1\omega$, and $|a|^{2}_{\mathrm{th}}/M=2$, with $M=500$ sites.
Given that $\varepsilon_{f}$ do not alter the behavior of the decay rate of $|a|^{2}$, it is natural to ask whether varying $\varepsilon_{f}$ has any significant effect as well on the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$, and on the signatures of the delayed KZM. We answer the first question in Figs. 6(a) and 6(b), where we show the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$, respectively, with $\tau_{q}$ for different values of $\varepsilon_{f}$. We can see that varying $\varepsilon_{f}$ does not significantly change the scaling of the KZM quantities considered. However, the $\tau_{q,c}^{*}$, shown as solid lines in Fig. 6(b), increases significantly as we decrease $\varepsilon_{f}$. This modification on $\tau_{q,c}^{*}$ becomes more apparent in Fig. 6(c), where we show the scaling of $\tau_{q,c}$ and $\tau_{q,c}^{*}$ as a function of $\varepsilon_{f}$. Notice that both quantities are inversely proportional to $\varepsilon_{f}$, with $\tau_{q,c}^{*}$ dropping faster than $\tau_{q,c}$ as $\varepsilon\rightarrow\infty$. As a result, the delayed KZM regime vanishes for large $\varepsilon_{f}$, highlighting that its signatures become more apparent for strongly dissipative systems quenched near criticality. We finally note that $\tau_{q,c}^{*}$ follows a power-law scaling as evidenced by the power-law fit curve shown in Fig. 6(c). In particular, since $\tau_{q,c}^{*}$ becomes the true critical quench time separating the breakdown and validity of the KZM at large $\varepsilon_{f}$, we expect that it should follow the power-law scaling [40]
@@
So far, we have shown that the presence of the delayed KZM leads to a significant deviation between the true freeze-out time, $\hat{t}$, and the transition time, $\hat{t}_{\mathrm{th}}$. Given that the delayed KZM becomes more prominent near criticality at strong dissipation, we now address in the next section how the threshold-based criterion for determining $\hat{t}_{\mathrm{th}}$ contributes to the deviation and whether a more accurate method can be used to measure $\hat{t}$.
@@ IV Transition Time Measurement
Figure 7: (a), (b) Scaling of (a) $\hat{t}_{\mathrm{th}}$ and (b) $\varepsilon(\hat{t}_{\mathrm{th}})$ of the open DM as a function of $\tau_{q}$ for different values of $|a|^{2}_{\mathrm{th}}$. (c) Boundary of the delayed KZM regime as a function of $\kappa$ for different $|a|^{2}$. The remaining parameters are set to $\varepsilon_{f}=0.2$ and $\kappa=0.1\omega$.
We answer the first question in Figs. 7(a) and 7(b), where we present the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$, respectively. For this part, while we only consider the open DM, the results here should apply to the COS and the open DLM as well. We can observe that the scaling of $\hat{t}_{\mathrm{th}}$ and $\varepsilon(\hat{t}_{\mathrm{th}})$ do not significantly change as we increase the threshold value. In particular, while the $\hat{t}_{\mathrm{th}}$ is only shifted by a constant value as $|a|^{2}_{\mathrm{th}}$ increases, both KZM quantities considered eventually collapse in a single scaling as $\tau_{q}\rightarrow\infty$. As for the boundaries of the delayed KZM regime, we can observe in Fig. 7(c) that the gap between $\tau_{q,c}$ and $\tau_{q,c}^{*}$ widens as we increase $|a|^{2}_{\mathrm{th}}$, implying that the delayed KZM becomes more prominent at large $|a|^{2}_{\mathrm{th}}$.
We can understand the widening of the delayed KZM regime for large $|a|^{2}_{\mathrm{th}}$ by noting that in an ideal setup where $\hat{t}$ can be accurately identified, the gap between $\tau_{q,c}$ and $\tau_{q,c}^{*}$ vanishes, and thus following the prediction in Ref. [40], $\tau_{q,c}=\tau_{q,c}^{*}=\hat{t}/\varepsilon_{f}$. Since for any threshold-based criterion, $\tau_{q,c}^{*}=\hat{t}_{\mathrm{th}}/\varepsilon_{f}$ and $\tau_{q,c}\neq\tau_{q,c}^{*}$ for large $\kappa$ and small $\varepsilon_{f}$, then
@@
Let us assume that within the time interval $[\hat{t},\epsilon_{f}\tau_{q}]$, the total occupation is exponentially growing such that $|a|^{2}\propto\exp(\gamma_{g}t)$, where $\gamma_{g}$ is the growth rate of the total occupation number. This assumption is supported by Fig. 4(c), where the $|a|^{2}$ of the open DM exponentially grows from the minimum value to its saturation value. With this assumption, we can infer that $\hat{t}_{\mathrm{th}}\propto\ln|a|^{2}_{\mathrm{th}}/\gamma_{g}$ and $\hat{t}\propto\ln|a|^{2}_{\mathrm{min}}/\gamma_{g}$, where $|a|^{2}_{\mathrm{min}}$ is the minimum value of the total occupation number. Thus,
@@
Figure 8: (a) Ramp protocol for the quench time within the $\tau_{q}$-regime of the delayed KZM for the open DLM. The circle point marks $\varepsilon(\hat{t})$, while the diamond point denotes $\varepsilon(\hat{t}_{\mathrm{th}})$. (b), (c) Spatiotemporal dynamics of (b) the occupation number $|a_{\ell}|^{2}$ and (c) the phase of the $\hat{a}_{\ell}$-mode, $\varphi_{\ell}$. (d) Exemplary dynamics of the defect number and the site-averaged $\varphi_{\ell}$. The vertical dashed lines correspond to $\omega\hat{t}$, while the solid lines represent $\omega\hat{t}_{\mathrm{th}}$. The remaining parameters are set to $\varepsilon_{f}=1.25$, $\kappa=J=0.1\omega$, and $|a|^{2}_{\mathrm{th}}=0.05|a|^{2}_{s}$, where $|a|^{2}_{s}$ is the steady state of $|a|^{2}$.
Now, determining an optimal threshold value that suppresses the signatures of the delayed KZM may be difficult to achieve as it requires prior knowledge of $|a|^{2}_{\mathrm{min}}$ for arbitrary $\tau_{q}$. This problem motivates the question of whether an alternative method can be used to infer $\hat{t}$ without relying on any threshold-based criterion. As we have hinted in the dynamics of the phase of the $\hat{a}$ mode of the open DM shown in Fig. 4(d), we can do this by choosing an appropriate order parameter that rapidly reaches its steady state upon the system entering a phase transition. In the case of the open DM, this order parameter corresponds to the boson mode’s phase, $\varphi$. We demonstrate this method further in Figs. 8(b) and 8(c), where we show that for nonzero dimensional systems, like the open DLM, we can use the phase information of the $\hat{a}_{\ell}$ modes to extract $\hat{t}$. As presented in Fig. 8(d), we can do this by determining the time at which either the defect number, $N_{d}$, or the site-averaged phase begins to saturate. In the COS level, the inferred $\hat{t}$ for this method would be equivalent to the moment the system picks a new global minimum it would fall onto, signaling phase transition. Thus, we expect that if the system’s phase information is available in an experimental setup, such as in Ref. [55], then that can serve as a more sensitive tool for detecting phase transitions compared to threshold-based order parameters that depend on the mode occupations.